Botox Facial Care had earnings after taxes of $282,000 in 20X1 with 200,000 shares of stock outstanding. The stock price was $48.80. In 20x2, earnings after taxes increased to $386,000 with the same 200,000 shares outstanding. The stock price was $59.00 a. Compute earnings per share and the P/E ratio for 20X1. (The P/E ratio equals the stock price divided by earnings per share.) (Do not round intermediate calculations. Round your final answers to 2 decimal places.) Farnings per share $ 141 34 61 times PIE ratio b. Compute eamings per share and the PIE ratio for 20X2. (Do not round intermediate calculations. Round your final answers to 2 decimal places.) 1 93 Earnings per share PIE ratio $ 30.57 times c. Why did the P/E ratio change? (Do not round Intemediate calculations. Input your answers as percents rounded to 2 decimal places.) The stack price increased by percent while FPS increased by percent

I only need help with part c. Parts a and b are correct.
Answer #1
C) Stock price increase = new price-old price/old price
Stock price increased = 59-48.80/48.80 = 21%
EPS increased by = new eps-old eps/old eps
EPS increased by = 1.93-1.41 / 1.41 = 37%
therefore the stock price increased by 21% while EPS increased by 37%